OpenStudy (anonymous):

money is borrowed at 13% simple interest. after one year, $1007.96 pays off the loan. how much was originally borrowed?

OpenStudy (jdoe0001):

\(\bf Balance= P(1+rt) \\ \quad \\ r=13\%\implies \cfrac{13}{100}\implies 0.13\qquad t=1\qquad Balance=1007.96 \\ \quad \\ P=\textit{original amount or Principal} \\ \quad \\ Balance= P(1+rt)\implies \cfrac{Balance}{(1+rt)}=P\)

OpenStudy (anonymous):

Thnk you very much!

OpenStudy (jdoe0001):

yw